# Modeling the dynamics of an oil drop driven by a surface acoustic wave in the underlying substrate

### v3

@@ Appendix C Computational implementation
The considered system, given by Eq. (29), contains two components $u=(\tilde{h},\tilde{\cal P})$ ($N=2$) and we use two equations, namely, Eqs. (29) and (30) for ($n=1)$, corresponding to the solution depending on a single spatial variable, $\tilde{x}$.
@@
  • Row $1$ ($i=1$) for Eq. (29)
    $d_{11}=1,\quad c_{12}=\tilde{h}^{3},\quad\gamma_{1}=-\tilde{\cal C}\frac{3{\cal S}}{8\tilde{K}_{z}^{4}}\psi(\tilde{x},\tilde{h})\Big{(}2\tilde{K}_{z}^{2}\tilde{h}^{2}-1+e^{2\tilde{K}_{z}\tilde{h}}(1-2\tilde{K}_{z}\tilde{h})\Big{)}\bigg{]}$ (66)
  • Row $2$ ($i=2$) for Eq. (30)
    $c_{21}=-1,\quad a_{21}=-{\rm Bo},\qquad a_{22}=1,\qquad f_{2}=\frac{{\cal S}\tilde{C}_{z}}{2\tilde{K}_{z}}\psi(\tilde{x},\tilde{h})$ (67)
At the domain ends, we apply Dirichlet boundary conditions $\tilde{h}=\tilde{h}_{p}$, where $\tilde{h}_{p}$ is the precursor film thickness, and $\partial\tilde{h}/\partial\tilde{x}=0$. In order to achieve convergent numerical simulations, we discretize the domain using $\Delta\tilde{x}=\tilde{h}_{p}$.

### v4

@@ Appendix C Computational implementation
[…]
@@
  • Row 11 ($i=1$) for Eq. (29)
    $)]d_{11}=1,\quad c_{12}=\tilde{h}^{3},\quad\gamma_{1}=-\tilde{\cal C}\frac{3{\cal S}}{8\tilde{K}_{z}^{4}}\psi(\tilde{x},\tilde{h})\Big(2\tilde{K}_{z}^{2}\tilde{h}^{2}-1+e^{2\tilde{K}_{z}\tilde{h}}(1-2\tilde{K}_{z}\tilde{h}$$\Big)\bigg$] (67)
  • Row 22 ($i=2$) for Eq. (30)
[…]
At the domain ends, we apply Dirichlet boundary conditions $\tilde{h}=\tilde{h}_{\rm p}$, where $\tilde{h}_{\rm p}$ is the precursor film thickness, and $\partial\tilde{h}/\partial\tilde{x}=0$. In order to achieve convergent numerical simulations, we discretize the domain using $\Delta\tilde{x}=\tilde{h}_{\rm p}$.